Changes between Initial Version and Version 1 of UC20170313


Ignore:
Timestamp:
Jul 30, 2021, 7:48:49 AM (10 months ago)
Author:
trac
Comment:

--

Legend:

Unmodified
Added
Removed
Modified
  • UC20170313

    v1 v1  
     1== APO 3.5-m Users Committee Phonecon, 3/13/17 ==
     2
     3Attending:  Nancy Chanover, Mark Klaene, Rachel Kuzio de Naray, Bill Ketzeback, Joanne Hughes, John Wisniewski, Russet !McMillan, Jason Lozo, Sarah Tuttle, John Bally, Mary Beth Kaiser, Anne Verbiscer, Jon Holtzman, Ben Williams, and Bruce Gillespie
     4
     5----
     6
     7=== User feedback and comments from institutional representatives: ===
     8
     9* Washington - Ben had nothing new to report.
     10* Georgia State - Rachel had nothing new to report.
     11* Seattle - Joanne had nothing new to report.
     12* NMSU - Jon had nothing new to report.
     13* JHU - Mary Beth had nothing new to report.
     14* Oklahoma - John W. said that his students recently had a good training run at APO.
     15* Colorado - John B. had nothing new to report.
     16* Wyoming - Chip had nothing new to report.
     17* Virginia - Anne mentioned that she now has IRAF scripts for ARCTIC, yet to be tested.
     18* NAPG - No report.
     19* Adler - Mark H. sent Nancy an e-mail requesting that more use be made of the wiki for instrument information (see later discussion).
     20
     21----
     22
     23=== Telescope and Instruments Report: ===
     24
     25Mark’s highlights included the mention that recent power-cycle tests of !TripleSpec were unsuccessful; staff from UVa will be coming to APO in May and/or June to work on the instrument and !TripleSpec is now officially offline. For ARCTIC, the site engineering staff is reconsidering design changes for rotating the diffuser, and the camera will now spend the next several months in a non-rotating-diffuser state.  Also, a crack in the diffuser necessitated replacing it with a spare.  Upgrades to ARCSAT’s ACP control software were made, as were some bug fixes.  ARCSAT is back to being fully operational, although !SurveyCam may need a connector repaired in the future.  Mark’s detailed report follows below.
     26
     27
     28''3.5-m Telescope and Instruments Highlights, 2/2/17 through 3/8/17''
     29
     300) Overview
     31
     32Observing time has been interrupted with winds, clouds, and dust but overall fairly productive.
     33
     341) Telescope - Telescope was operational with only minimal issues
     35
     362) Instruments
     37
     38DIS blue and red cameras have been operational.  The red camera shows some indication of return of contamination but still at a minimal amount.  We continue to monitor frequently and have temperature humidity readings inside instrument as well.
     39
     40!TripleSpec was warmed and re-cooled in order to try address corrupted image issue.  The problem did not correct with thermal cycle and the instrument has been warmed and taken out of service until UVa team can address this on site.  Possibly in May.
     41
     42Agile has been operational this period.  A problem with rotator showed up and is a known infrequent communication issue with the new TCC protocol.  This is being looked into.
     43
     44Echelle has been operational with no problems.  We are looking for time in the spring to perform vacuum servicing.
     45
     46NIC-FPS is operational.  A lot of testing with the Grism has been going on and documentation is in work to help supplement the users manual on Grism use.
     47
     48ARCTIC is operational. The diffuser rotation has been problematic.  We are designing a new holder so that the diffuser can be installed while the rotational issues can be worked on the bench.  An SDSS g filter was damaged and developed a chip on the very edge.  Work is ongoing to seal the chip.  It does not appear to effect the optical qualities in beam.
     49
     503) ARCSAT - Major ACP software upgrade was done and that hopes to fix some issues.  However, it created a lot of broken links and so not all previous functionality is back yet.  These are being worked on around 3.5m priorities.  The telescope is operational with these remaining issues.
     51
     52----
     53
     54=== Report on DIS upgrades PDR ===
     55
     56Nancy said that the DIS upgrade PDR final report will be distributed shortly to all users.  Sarah added that the PDR went well, and helped facilitate in the gathering of a collection of DIS engineering documentation from the past 10+ years.  As a result of the PDR, there are various tests and upgrades being planned for DIS over the summer, albeit nothing major, yet.  Russet is helping to track the environmental issues that are affecting the instrument. Lastly, decisions will need to be taken on what DIS upgrades to implement in light of the new spectrograph project.
     57
     58----
     59
     60=== Q1 3.5m schedule updates ===
     61
     62Russet noted that all Q1 open time has been assigned.
     63
     64----
     65
     66=== Q2 3.5m scheduling ===
     67
     68Ben announced that the Q2 schedule was just posted, and he’s received no complaints so far.  The revised cover sheet for the proposals proved to be a great success and was helpful; it will be used for future scheduling quarters.  Some of the April dark time was oversubscribed, and there is some residual DD dark time left in June. Lastly, the ARCSAT Q2 proposals are due next week.
     69
     70----
     71
     72=== Update on ARCTIC diffuser ===
     73
     74Nancy said that everyone should be aware that the ARCTIC diffuser is available to all users — it is an asset for observing bright targets, even without diffuser rotation.  John W. added that the diffuser enables high-precision photometry that is better than the classic method of defocussing, and with the diffuser, ARCTIC photometric precision rivals that of Kepler.  Mark said that while we consider modifications to the diffuser rotation system, it would be very useful if we knew if users had any performance requirements, such as for rotation speed and stability — send any thoughts on this to Mark or Leslie Hebb.
     75
     76
     77----
     78
     79=== Preliminary schedule for summer shut down, DIS upgrades ===
     80
     81Mark said that the tentative shut down dates are for a 31 July start with a (weather-permitting) restart of science observing on 18 August.  The DIS engineering down time will be 4 through 12 September (starting 2nd halves, respectively), and is also a tentative schedule at this time.  Also, the echelle vacuum will be under routine maintenance for 7 to 10 days at the end of April.
     82
     83
     84----
     85
     86
     87=== Encouraging use of APO 3.5m wiki ===
     88
     89Nancy pointed to the APO 3.5m wiki, which is hosted by NMSU at http://astronomy.nmsu.edu:8000/apo-wiki/wiki/WikiStart.  This site seems to be the natural place to keep and maintain our instrumentation and other documents, where they can be easily edited.  Nancy would like everyone to make more use of the wiki, and she will start placing User Committee documentation there.
     90
     91
     92----
     93
     94
     95=== ACTION ITEMS:===
     96
     97'''open from previous meetings:'''
     98
     99* Nancy will work with Jason, Rachel, Jon H. and others to get the ARCTIC information on the APO wiki cleaned up and broadly accessible to ARC users.
     100Status:  Closed, Nancy will send e-mail to the 3.5m general mailer.
     101
     102
     103'''new from this meeting:'''
     104None
     105
     106
     107----
     108
     109=== Next meeting:===
     110The next Users Committee phonecon will be on Monday 10 April at 1030 MDT — the agenda and other materials will be sent to committee members during the preceding week.